OpenStudy (anonymous):

Identify the 19th term of the arithmetic sequence 5, 16, 27 …

Parth (parthkohli):

Use the given formula:\[\rm a_n = a_1 + (n - 1)d \]

OpenStudy (anonymous):

Yes i am having trouble what n would be

OpenStudy (anonymous):

a19=5+(n-1)11

Parth (parthkohli):

They are asking for 19th term, so \(\rm n = 19\).

Parth (parthkohli):

\[\rm a_{19} = 5 + (19 - 1)\times 11\]

Parth (parthkohli):

\[\rm = 5 + (18\times 11)\]

OpenStudy (anonymous):

5+(198)

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

OH okay i see. So n and a= are the same?

Parth (parthkohli):

“a” is just the name of that sequence.

Parth (parthkohli):

So, \(\rm a_{19}\) reads, “nineteenth term of the sequence \(\rm a\)”

Parth (parthkohli):

You could name the sequence anything. You could name it \(\rm s\) too.

Parth (parthkohli):

But people mostly use \(\rm a\) for sequences. lol

OpenStudy (anonymous):

oh okay thank you!!

Parth (parthkohli):

You're welcome!

Parth (parthkohli):

ParthKohli*

Parth (parthkohli):

lol